OpenStudy (anonymous):

Which equation is not equivalent to the formula m = ca? c equals m over a a equals c over m a equals m over c m = ac @iGreen. im stuck

OpenStudy (anonymous):

D you can cross out

OpenStudy (anonymous):

D is equivalent to m = ca.. 'ca' and 'ac' are the same thing..

OpenStudy (anonymous):

So D isn't your answer.

OpenStudy (anonymous):

i know i said you can cross it out

OpenStudy (anonymous):

You can solve this equation 2 ways. By dividing 'c' or 'a' to both sides. Let's do 'c'. Divide 'c' to both sides: \(m = ca\) \(\dfrac{m}{c} = a\) So we know that a = m/c..so what's your answer going to be?

OpenStudy (anonymous):

c

OpenStudy (anonymous):

No..a = m/c, that means it's EQUIVALENT to it, and we're looking for one that ISN'T equivalent.

OpenStudy (anonymous):

ohhh so uhm i tihnk its B

OpenStudy (anonymous):

Yep, you got it. Since \(a = \dfrac{m}{c}\), it can't equal \(\sf \dfrac{c}{m}\).